Solve by substitution method and what is the solution set?
0.1x+0.8=-4.6
0.9x-0.5y=4.8

Answer by mananth(16946)

0.1 x + 0.8 y = -4.6 .............1
-0.5 y = 4.8 -0.9 x ---------2
/ -0.5 =
y = -9.6 + 1.8 x
Substitute y in (1)
0.1 x + 0.8 ( -9.6 + 1.8 x ) = -4.6
0.1 x -7.68 + 1.44 x = -4.6
0.1 x + 1.44 x = -4.6 7.68
1.54 x = 3.08
/ 1.54
x= 2
Plug the value of x in (1)
0.1 x + 0.8 y = -4.6
0.1 * 2 + 0.8 y = -4.6
0.2 + 0.8 y = -4.6
0.8 y = -4.8
/ 0.8
y= -6